# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ Abstract
This review brings together a body of experimental, theoretical, and simulation work addressing the behaviour of high-field devices and the fundamental process of vacuum arcing, also referred to as vacuum breakdown. Detailed experimental observations and results relating to a wide range of aspects of high-field devices including conditioning, field and temperature dependence of breakdown rate, and the field-holding capability as a function of device geometry and of different materials are presented. The different observations are then addressed theoretically, and with simulation, capturing the sequence of processes that lead to vacuum breakdown and explaining the major observed experimental dependencies. The core of the work described in this review was carried out by a broad multi-disciplinary collaboration in an over a decade-long program to develop high-gradient, 100 MV/m-range, accelerating structures for the CLIC project, a possible future linear-collider high-energy physics facility. Connections are made to the broader linear collider, high-field, and breakdown communities.
@@ I.1 General introduction to breakdown and context for this work
Although vacuum breakdowns play a critical role in a wide variety of applications, the fundamentals of their initiation have remained an open question. However, the knowledge of these mechanisms may dramatically improve the controllability of both desired and failure mode-type systems. In general terms, a system starts with an applied high electrostatic or radio-frequency electromagnetic field, possibly producing low-level currents caused by electron field emission. Field emission currents are typically a micro-ampere or less. This metastable high-field state can be sustained over long periods before the system undergoes an abrupt, nanosecond timescale, transition to a breakdown. Systems for which breakdown is a failure mode are designed to operate in the metastable insulating state. When a breakdown occurs at some position on a high-field surface, the insulating vacuum becomes conductive through the formation of a plasma that carries high current, 100s of amperes and more, between electrodes or through the volume of radio frequency structures. The accelerated high current draws power from the power source that sets up the high field and causes it to drop to a small fraction of that in the insulating state. This field collapse constitutes a failure in some systems, while the high current and lower field is the operating mode in others. This report focuses on systems of the former type.
@@
As we will see in this review, breakdowns occur randomly in both space and time. Breakdown nucleation sites are very small, some tens of nanometer-scale or less, the upper limit coming from observations, while device surfaces typically have dimensions from millimeters even up to meters. The onset of breakdown occurs in nanoseconds while fields, sometimes continuous and sometimes pulsed, can be applied for hours, days, or even longer depending on the application. The randomness and wide range of time and length scales make measurements and analyses of the processes underlying the initiation and early-stage evolution of breakdowns challenging. The work described in this report addresses this challenge.
@@
Vacuum breakdown has been the focus of dedicated studies by different fields affected by it, or using it, for over a century. Several books are devoted to summarizing the experimental evidence, the early modelling work and the technological consequences and uses of vacuum arcs [88, 93, 112, 94, 18, 5]. The arc initiation dynamics in particular has been the focus of several reviews [30, 70] drawing from seminal works dedicated to the study and consequences of field-emitted electrons on the initiation of vacuum breakdowns [71, 103, 175, 24, 180]. The elusiveness of breakdown initiation has nevertheless prevented a full understanding of the underlying mechanisms, in particular establishing a detailed physical origin of the observed breakdown rates.
@@
In recent years there has been a major new effort to understand the fundamentals of vacuum arcs driven this time by a push to develop TeV-range energy electron-positron linear colliders for high-energy physics. There have been several different projects to design and develop key technology for such a collider including CLIC and NLC/JLC [2, 91]. To limit length, and consequently cost, linear colliders depend on high-gradient accelerating structures. The rf (radio frequency) accelerating structures for CLIC target an accelerating gradient of 100 MV/m, resulting in surface electric fields over 200 MV/m. This high field makes vacuum breakdown a major limitation.
@@
The limitation of gradient by breakdown gives a clear scientific goal and has motivated a long-term effort to deepen the understanding of breakdown in order to achieve as high an accelerating gradient as possible, while minimizing the breakdown rate. This review will summarize the latest experimental and corresponding theoretical work carried out towards understanding vacuum breakdown, with a particular focus on the processes that lead to its initiation. The review draws mainly on work carried out by a multidisciplinary collaboration loosely centered on the CLIC study but also includes important related work carried out by other groups, within and outside of linear colliders studies.
@@
In addition, the progress described in this report has been made possible also due to advances in development of simulation methodologies and tools, which had not been earlier either available or applied to address the physics underlying the vacuum arcing. Generally, to develop a consistent theoretical concept explaining the onset of vacuum arcing, the construction of a single multiscale and multiphysics simulation model where all relevant processes are interconnected is necessary. The processes to be included are the distribution of electron densities at surface irregularities as a result of applied electric fields, modified by the non-uniform distribution of surface charge atomic dynamics and the subsequent effect of Maxwell stress on the defect dynamics under the surface along with the complex physics of interaction of field emission currents with metal tips resulting in an eventual plasma buildup. Simulation of plasma discharge can enable the link between the theory and experimental observations of post-breakdown surface features.
@@
Until recently the theoretical efforts in this direction were limited. The effects of electric field on materials, for example, were considered either within analytical theories developed for field emission applications [51, 50] or within static quantum-mechanical calculations, see e.g. [104, 15, 153, 152]. The dynamic effects in materials in the field emission regime were addressed in the studies of thermal response of intensively field emitting microtips, however, within a static tip approximation [55]. Lately, developments of atom probe tomography [116, 179] based on field-evaporation phenomena [49] triggered a number of density functional theory calculations, which provided interesting insights on electron density distributions near metal surfaces in the presence of high electric fields [149]. All these approaches were developed to address specific technological problems and have not provided a consistent theoretical picture of physics triggered in metals exposed to high electric fields. Neither the atom-level insights on the modification of surface morphology or defect dynamics within the bulk of the materials exposed to high electric fields were available.
@@
First attempts to apply molecular dynamic methods to the vacuum arcing phenomenon date back to the beginning of 2000s, where the surface atoms were modeled as charged ions to which the pulling force due to an applied electric field (a fixed maximal value) was applied [128, 59]. The lack of physical background of the proposed model limited the application of this approach to explain the vast majority of experimental evidence collected in many laboratories. Later, more rigorous approaches [36, 37, 172, 174] were proposed that were developed to follow the changes within the metal exposed to a high electric field. These simulations have enabled many non-intuitive insights into the processes governing the onset of vacuum breakdowns.
@@ I.2 An overview of vacuum breakdown
We begin with a high surface electric field which applies a tensile stress to a metal (conducting) surface. This tensile stress leads to dislocation activity within the bulk metal of the electrodes. This activity, also referred to as plastic response, subsequently leads to the formation of the surface structures where breakdowns are initiated. Plastic response occurs at levels well below the stress limit of the material, in the form of dynamics of point-like defects. Specifically it drives the creation, movement and annihilation of dislocations. Due to energy minimization, dislocations project to the electrode surface creating the surface features that will eventually evolve to breakdown sites. Dislocation dynamics are the origin of statistical nature of the experimentally observed dependencies of breakdown rate (the number of breakdowns divided by the number of applied field pulses over a long operation period) on field level, temperature and material. They are also explain an important part of the conditioning process, by which the voltage holding capability can be improved by repeated or extended application of high voltage.
@@
On the surface of electrodes, in particular on the cathode, several mechanisms occur when a high electric field is applied. In the discussion of surface dynamics, it is useful to distinguish between intrinsic and extrinsic mechanisms. Intrinsic mechanisms are related to the dynamics of material from which the electrodes are made, for example copper in most cases in this report. Extrinsic mechanisms are related to materials other than that of the bulk of the electrodes, for example oxides, carbon or dust-like contaminants. In general terms extrinsic mechanisms dominate the early stages of operation of a high field system while intrinsic mechanisms dominate later stage behavior and determine the ultimate operating conditions, as the high field limit of the surface is approached late in the conditioning process.
@@
The most important intrinsic dynamic surface processes driven by high surface electric fields include the biased diffusion of surface atoms, localized stress due to local field enhancement and heating due to field emission. Biased diffusion is the migration of surface atoms in the direction of increasing surface field. Biased migration occurs because the surface electric field polarizes surface atoms and consequently have an attractive force towards an increasing field. This means that surface atoms preferentially migrate toward areas of locally higher electric field. Such a higher local field occurs when there are surface features so the biased diffusion gives a feedback mechanism to sharpen an electrode surface topography. The local field enhancement gives locally enhanced electron field emission, which results in localized heating. The geometrical field enhancement and subsequent heating due to field emission result in increased local stresses, which contribute further but locally to the plastic mechanisms
@@
The next aspects of the breakdown process are the mechanisms through which the vacuum is filled with neutral and charged particles, which give rise to the insulating-to-conducting transition that causes applied fields to collapse, the most directly observable aspect of breakdown. Here we also add electron field emission, a low-level current that occurs at sufficiently high fields, even when a breakdown does not occur. When a surface feature exceeds a certain threshold, either through growth driven by underlying surface dynamics projecting to the surface or surface dynamics, localized field-emission driven heating drives atomic evaporation. Field emitted, and then accelerated currents, as well as the applied field itself ionize atoms, creating a back-bombardment on the cathode surface, liberating further atoms. This then gives rise to the avalanche-like process that forms the current carrying plasma above the cathode surface.
@@ II.1 Overview and context
In this section, we describe the main experimental systems that were used to make the high-field and breakdown measurements that provide the foundation for the analysis reviewed here. There are of course a wide range of systems in which breakdown occurs, as summarized in the introduction, but generally there is very little quantitative data available from them. This lack is due to various reasons that include operation of devices limited to a narrow range of specified parameters, lack of instrumentation, lack of systematic measurements, concern about damaging equipment and protection of intellectual property (which is often the case in commercial devices). The detailed study and construction of a theoretical understanding of a phenomenon as complicated as vacuum breakdown, however requires high quality measurements of relevant physical quantities. Dependencies and repeatability must be determined as well as the range of applicability. Ideally equivalent data should be taken by different groups in different setups.
@@
Fortunately in recent years such a body of measurements has become available thanks to extensive experimental programs carried out as part of the development of high-energy linear colliders, as well as smaller-scale projects that have common high-field holding requirements. The linear collider programs have provided the motivation, but the results and understanding are important for the broader high-field community.
@@
Linear colliders are a class of possible future high-energy physics facilities that aim for electron positron collisions with center of mass energies in 250 GeV extending into the TeV range. The electron positron linear collider project with the highest targeted collision energy is CLIC, which foresees an initial stage with 380 GeV center of mass collisions, and extensions to energies up to 3 TeV. In addition to CLIC, other linear collider projects include the NLC/JLC which was formally ended in 2004, C3 and the ILC, International Linear Collider. The latter is based on superconducting cavities so has very different technological challenges and is not directly relevant for this report. References to the NLC/JLC and CLIC projects can be found here NLC/JLC [91] and CLIC [2].
@@
The very high collision energy needed for a linear collider motivates a push towards a very high accelerating gradient in order to minimize the length, and consequently the cost of a facility. Taking CLIC as an example, the design accelerating gradient is 72 MV/m for the initial 380 GeV stage and 100 MV/m for the higher energy stages up to 3 TeV. This results in over 5 km of active accelerating linac length at 380 GeV, with a corresponding total facility length of 11 km, and over 30 km active length at 3 TeV.
@@
Vacuum breakdown is a major limitation in achieving such accelerating gradients. This can be seen by considering the surface electric fields that occur inside the accelerating structures. The geometry of linear collider structures is such that the peak surface electric fields are approximately 2 to 2.5 times higher than the accelerating gradient. Thus for accelerating gradients the range of 70 to 100 MV/m peak surface electric fields are in excess of 200 MV/m. This is well into the range of surface electric field where vacuum breakdown can occur. Developing accelerating structures that can operate reliably at such high surface electric field, and over the multi-decade lifetime of high energy facilities, has been a major challenge.
@@
Given the importance of a high accelerating gradient for CLIC, the project has invested in prototype accelerating structure development program coupled with a multi-disciplinary study into the fundamental processes of breakdown. The main objective of the prototype development program has been to validate the feasibility and practicality of operating accelerating gradients at the design values of 70 to 100 MV. The main objectives the study into fundamental processes has been to improve performance through understanding, including increasing achievable gradient through radio frequency design, through material and process choices, optimization of the overall facility (in power consumption and cost), reduction in conditioning time and optimization of operational issues such as recovery from breakdown. All these can be better addressed through an improved understanding of the fundamental processes of breakdown.
@@
To carry out the prototype development program and the fundamental studies, the CLIC project has made a major investment in experimental infrastructure, in particular in test stands where high-field experiments can be carried out in controlled and well instrumented manner. They consist of radio frequency test stands, driven by klystrons operating in X-band and referred to as the XBoxes, 11.994 GHz, an S-band, 2.9985 GHz, and pulsed dc test stands, driven by Marx generators. The radio frequency test stands are capable of powering accelerating structures to full specifications but are rather costly, as are prototype accelerating structures. The pulsed dc systems have been designed to allow directly comparable experiments to those carried out in the radio frequency test stands but using an apparatus that is much simpler, cheaper and more easily instrumented and make tests on relatively simple electrodes. There is an overlap, or near overlap in some cases, in important parameters while an even even larger range of certain parameters, for example repetition rate. Because of the overlap, protocols can be reproduced and results can be directly, and quantitatively compared.

### v2

# Fundamentals of Vacuum Breakdown in High-Field Systems

@@ Abstract
This review consolidates experimental, theoretical, and simulation work examining the behavior of high-field devices and the fundamental process of vacuum arc initiation, commonly referred to as breakdown. Detailed experimental observations and results relating to a wide range of aspects of high-field devices, including conditioning, field and temperature dependence of breakdown rate, and the ability to sustain high electric fields as a function of device geometry and materials, are presented. The different observations are then addressed theoretically, and with simulation, capturing the sequence of processes that lead to vacuum breakdown and explaining the major observed experimental dependencies. The core of the work described in this review was carried out by a broad multi-disciplinary collaboration in an over a decade-long program to develop high-gradient, 100 MV/m-range, accelerating structures for the CLIC project, a possible future linear-collider high-energy physics facility. Connections are made to the broader linear collider, high-field, and breakdown communities.
@@ I.1 General introduction to breakdown and context for this work
Although vacuum breakdowns play a critical role in a wide variety of applications, the fundamentals of their initiation have remained an open question. However, the knowledge of these mechanisms may dramatically improve the controllability of both desired and failure mode-type systems. In general terms, a device starts with an applied high electrostatic or radio-frequency electromagnetic field, possibly producing low-level currents caused by field electron emission, i.e. emission of electrons due to extremely high surface electric fields, frequently referred to also as field emission. Field emitted currents are typically a micro-ampere or less. This metastable high-field state can be sustained over long periods before the device undergoes an abrupt, nanosecond timescale, transition to a breakdown. Systems for which breakdown is a failure mode are designed to operate in the metastable insulating state. When a breakdown occurs at some position on a high-field surface, the insulating vacuum becomes conductive through the formation of a plasma that carries high current, 100s of amperes and more, between electrodes or through the volume of radio frequency structures. The accelerated high current draws power from the power source that sets up the high field and causes it to drop to a small fraction of that in the insulating state. This field collapse constitutes a failure in some systems, while the high current and lower field is the operating mode in others. This report focuses on systems of the former type.
@@
As we will see in this review, breakdowns occur randomly in both space and time. Breakdown nucleation sites are very small, some tens of nanometer-scale or less, the upper limit coming from observations, while device surfaces typically have dimensions from millimeters even up to meters. The onset of breakdown occurs in nanoseconds while fields, sometimes continuous and sometimes pulsed, can be applied for hours, days, or even longer depending on the application. The combination of randomness and wide span of time and length scales make it very challenging to measure and analyse the processes underlying the initiation and early-stage evolution of breakdowns. The work described in this report addresses this challenge.
@@
Vacuum breakdown has been the focus of dedicated studies by different fields affected by it, or using it, for over a century. Several books are devoted to summarizing the experimental evidence, the early modelling work and the technological consequences and uses of vacuum arcs [93, 98, 121, 99, 19, 5]. The arc initiation dynamics in particular has been the focus of several reviews [33, 74] drawing from seminal works dedicated to the study and consequences of field-emitted electrons on the initiation of vacuum breakdowns [75, 110, 185, 26, 190]. The elusiveness of breakdown initiation has hindered a full understanding of the underlying mechanisms, particularly the physical origins of the observed breakdown rates.
@@
In recent years there has been a major new effort to understand the fundamentals of vacuum arcs, driven this time by a push to develop TeV-range energy electron-positron linear colliders for high-energy physics. There have been several different projects to design and develop key technology for such a collider including CLIC and NLC/JLC [2, 96]. To limit length, and consequently cost, linear colliders depend on high-gradient accelerating structures. The rf (radio frequency) accelerating structures for CLIC target an accelerating gradient of 100 MV/m, resulting in surface electric fields over 200 MV/m. This high field makes vacuum breakdown a major limitation.
@@
The limitation of gradient by breakdown gives a clear scientific goal and has motivated a long-term effort to deepen the understanding of breakdown in order to achieve as high an accelerating gradient as possible, while minimizing the breakdown rate, i.e. the frequency at which breakdowns occur. This review will summarize the latest experimental and corresponding theoretical work carried out towards understanding vacuum breakdown, with a particular focus on the processes that lead to its initiation. The review draws mainly on work carried out by a multidisciplinary collaboration loosely centered on the CLIC study but also includes important related work carried out by other groups, within and outside of linear colliders studies.
@@
In addition, the progress described in this report has been made possible due to advances in development of simulation tools and methodologies, which had not been earlier either available or applied to address the physics underlying vacuum breakdown. Generally, to develop a consistent theoretical concept explaining the onset of vacuum arcs, the construction of a single multiscale and multiphysics simulation model where all relevant processes are interconnected is necessary. The processes to include span (i) the redistribution of electron densities on the metal surface as a result of applied electric fields, (ii) the consequent non-uniform distribution of field-induced charges especially on surface irregularities, (iii) the modification of atomic dynamics due to the resulting field-induced stress, especially on surface and under-surface defects, and (iv) the complex interaction of field emission currents with metal tips resulting in their thermal runaway, evaporation, and an eventual plasma buildup. The simulation of plasma initiation can enable the link between the theory and experimental observations of post-breakdown surface features.
@@
Until recently the theoretical efforts in this direction were limited. The effects of electric field on materials, for example, were considered either within analytical theories developed for field emission applications [54, 53] or within static quantum-mechanical calculations, see e.g. [111, 16, 162, 161]. The dynamic effects in materials in the field emission regime were addressed in the studies of thermal response of intensively field emitting microtips, however, within a static tip approximation [58]. Lately, developments of atom probe tomography [125, 189] based on field-evaporation phenomena [52] triggered a number of density functional theory calculations, which provided interesting insights on electron density distributions near metal surfaces in the presence of high electric fields [158]. All these approaches were developed to address specific technological problems but have not provided a consistent theoretical picture of physics triggered in metals exposed to high electric fields. Neither the atom-level insights on the modification of surface morphology or defect dynamics within the bulk of the materials exposed to high electric fields were available.
@@
First attempts to apply molecular dynamic methods to the vacuum breakdown phenomenon date back to the beginning of 2000s, where the surface atoms were modeled as charged ions to which the pulling force due to an applied electric field (a fixed maximal value) was applied [137, 62]. The lack of physical background of the proposed model limited the application of this approach to explain the vast majority of experimental evidence collected in many laboratories. Later, more rigorous approaches [39, 40, 182, 184] were proposed that were developed to follow the changes within the metal exposed to a high electric field. These simulations have enabled many non-intuitive insights into the processes governing the onset of vacuum breakdowns.
@@ I.2 An overview of vacuum breakdown
A schematic representation of the main stages of the initiation of an arc is shown in Fig. 1. In general terms, we are considering a system with a starting configuration of a metal electrode or radio frequency structure in vacuum to which electromagnetic fields are applied. The types of structures that are considered in this report are made from copper and typically have some tens of nanometer level surface finish. The fields act on the entire structure surface setting in motion processes, dislocation dynamics. in the material and on its surface, surface diffusion, that occur everywhere that fields are high. The consequence of these processes is the formation of small, 10s of nm scale, features which enhance local fields. The local field enhancement induces further processes, the emission of electrons and neutral copper atoms. Occasionally, one of these features falls into an avalanche condition, which results in the formation of a conductive plasma that effectively short-circuits the system. This is what is observed as a breakdown. These breakdown initiation stages are relevant for a highly conditioned structure, that is one that has been exposed extensively to high fields. The mechanisms are intrinsic so are determined, and limited, by the characteristics of the material. In lower-field devices, breakdown initiation may be dominated by contaminants, that is, extrinsic mechanisms. The different stages with distinct physical processes have been identified and are represented by the different boxes in Fig. 1. We continue by elaborating these stages briefly in the rest of this section. The different stages are discussed in depth in sections IV to IX.
@@
Surface electric fields apply a tensile stress to a metal (conducting) surface, that extends into the bulk metal of the electrodes. This tensile stress leads to dislocation activity within the bulk. This activity, also referred to as plastic response, subsequently leads to the formation of the surface structures where breakdowns are initiated. Plastic response occurs at levels well below the stress limit of the material, in the form of dynamics of point-like defects. Specifically, it drives the creation, movement, and annihilation of dislocations. Due to energy minimization, dislocations project to the electrode surface, creating the surface features that will eventually evolve to breakdown sites. Dislocation dynamics are the origin of the statistical nature of the experimentally observed dependencies of breakdown rate (the number of breakdowns divided by the number of applied field pulses over a long operation period) on field level, temperature and material. They also explain an important part of the conditioning process, by which the field holding capability, i.e. the maximum field a system can hold without breakdown, can be improved by repeated or extended application of high voltage.
@@
On the surface of electrodes, in particular on the cathode, several mechanisms occur when a high electric field is applied. In the discussion of surface dynamics, it is useful to distinguish between intrinsic and extrinsic mechanisms. Intrinsic mechanisms are related to the dynamics of material from which the electrodes are made, for example copper in most cases in this report. Extrinsic mechanisms are related to materials other than that of the bulk of the electrodes, for example oxides, carbon or dust-like contaminants. In general terms, extrinsic mechanisms dominate the early stages of operation of a high field system while intrinsic mechanisms dominate later stage behavior and determine the ultimate operating conditions, as the high field limit of the surface is approached late in the conditioning process.
@@
The most important intrinsic dynamic surface processes driven by high surface electric fields include the biased diffusion of surface atoms, localized stress due to local field enhancement and heating due to field emission. Biased diffusion is the migration of surface atoms in the direction of increasing surface field. Biased migration occurs because the surface electric field polarizes surface atoms and consequently induces an attractive force towards an increasing field. This means that surface atoms preferentially diffuse toward areas of locally higher electric field. Such a higher local field occurs when there are surface features so the biased diffusion gives a feedback mechanism to sharpen an electrode surface topography. The local field enhancement gives locally enhanced electron field emission, which results in localized heating. The geometrical field enhancement and subsequent heating due to field emission result in increased local stresses, which contribute further but locally to the plastic mechanisms.
@@
The next aspects of the breakdown process are the mechanisms through which the vacuum is filled with neutral and charged particles, which give rise to the insulating-to-conducting transition that causes applied fields to collapse, the most directly observable aspect of breakdown. Here we also add electron field emission, a low-level current that occurs at sufficiently high fields, even when a breakdown does not occur. When a surface feature exceeds a certain threshold, either through growth driven by underlying surface dynamics projecting to the surface or surface dynamics, localized field-emission driven heating drives atomic evaporation. Field emitted, and then accelerated electrons, as well as the applied field itself ionize atoms, creating a back-bombardment on the cathode surface, liberating further atoms. This then gives rise to the avalanche-like process that forms the current-carrying plasma above the cathode surface.
@@
Finally, table 1 in the Appendix gives a comprehensive list of the specialized terms, abbreviations, and symbols used throughout the paper for easy reference.
@@ II.1 Overview and context
In this section, we introduce the application, linear colliders, that has motivated this in-depth study of breakdown, and then describe the main experimental systems that have provided the fundamental results are the input for the development of the theoretical models described in this review. The requirements of linear colliders motivate the study of breakdown under very high surface fields.
@@
There are a wide range of systems in which breakdown occurs, as summarized in the introduction, but generally, there is very little quantitative data available from them. This lack is due to various reasons that include operation of devices limited to a narrow range of specified parameters, lack of instrumentation, lack of systematic measurements, concern about damaging equipment and protection of intellectual property (which is often the case in commercial devices). The detailed study and construction of a theoretical understanding of a phenomenon as complicated as vacuum breakdown, however, requires high-quality measurements of relevant physical quantities. Dependencies and repeatability must be determined as well as the range of applicability. Ideally, equivalent data should be taken by different groups in different setups.
@@
Fortunately, in recent years, such a body of measurements has become available thanks to extensive experimental programs carried out as part of the development of high-energy linear colliders, as well as smaller-scale projects that have common high-field holding requirements. The linear collider programs have provided the motivation, but the results and understanding are important for the broader high-field community.
@@
Linear colliders are a class of possible future high-energy physics facilities that aim for electron-positron collisions with center of mass energies in 250 GeV, extending into the TeV range. The electron positron linear collider project with the highest targeted collision energy is CLIC, which foresees an initial stage with 380 GeV center of mass collisions, and extensions to energies up to 3 TeV. In addition to CLIC, other linear collider projects include the NLC/JLC which was formally ended in 2004, C3 and the ILC, International Linear Collider. The latter is based on superconducting cavities, so it has very different technological challenges and is not directly relevant for this report. References to the NLC/JLC and CLIC projects can be found here NLC/JLC [96] and CLIC [2].
@@
The very high collision energy needed for a linear collider motivates a push towards a very high accelerating gradient to minimize the length, and consequently the cost of a facility. Taking CLIC as an example, the design accelerating gradient is 72 MV/m for the initial 380 GeV stage and 100 MV/m for the higher energy stages up to 3 TeV. This results in over 5 km of active accelerating linac length at 380 GeV, with a corresponding total facility length of 11 km, and over 30 km active length at 3 TeV.
@@
Vacuum breakdown is a major limitation to achieving such accelerating gradients. This can be seen by considering the surface electric fields that occur inside the accelerating structures. The geometry of linear collider structures is such that the peak surface electric fields are approximately 2 to 2.5 times higher than the accelerating gradient. Thus for accelerating gradients the range of 70 to 100 MV/m peak surface electric fields are in excess of 200 MV/m. This is well into the range of surface electric field where vacuum breakdown can occur. In addition, accelerating structures operating at such high gradients can also be subject to progressive deterioration due to pulsed surface heating, [100]. Developing accelerating structures that can operate reliably at such high surface electric field, and over the multi-decade lifetime of high energy facilities, has been a major challenge.
@@
Given the importance of a high accelerating gradient for CLIC, the project has invested in prototype accelerating structure development program coupled with a multi-disciplinary study into the fundamental processes of breakdown. The main objective of the prototype development program has been to validate the feasibility and practicality of operating accelerating gradients at the design values of 70 to 100 MV. The main objectives of the study into fundamental processes has been to improve performance through understanding, including increasing achievable gradient through radio frequency design, through material and process choices, optimization of the overall facility (in power consumption and cost), reduction in conditioning time and optimization of operational issues such as recovery from breakdown. All these can be better addressed through an improved understanding of the fundamental processes of breakdown.
@@
To carry out the prototype development program and the fundamental studies, the CLIC project has made a major investment in experimental infrastructure, in particular in test stands where high-field experiments can be carried out in controlled and well instrumented manner. They consist of radio frequency test stands referred to as the XBoxes, driven by klystrons operating at 11.994 GHz and 2.9985 GHz, and pulsed dc test stands, driven by Marx generators. The radio frequency test stands are capable of powering accelerating structures to full specifications but are rather costly, as are prototype accelerating structures. The pulsed dc systems have been designed to allow directly comparable experiments to those carried out in the radio frequency test stands but using an apparatus that is much simpler, cheaper and more easily instrumented and make tests on relatively simple electrodes. There is an overlap, or near overlap in some cases, in important parameters while an even larger range of certain parameters, for example repetition rate. Because of the overlap, protocols can be reproduced and results can be directly, and quantitatively compared.